OpenStudy (anonymous):

Solve for r. 5r − 3(2r + 5) = 9

OpenStudy (anonymous):

r=6

OpenStudy (anonymous):

show work

OpenStudy (lalaly):

5r-6r+15=9 -r=-6 r=6

OpenStudy (anonymous):

5r- 3(2r+5)=9 5r-6r + 15=9 5r-6r = 9-15 -r=-6 r=6

OpenStudy (anonymous):

5r - 3(2r + 5) = 9 5r - 3*2r - 3*5 = 9 5r - 6r - 15 = 9 -r = 9 + 15 -r = 24 r=-24

OpenStudy (anonymous):

coolnixy is correct